# Gravity and fluid dynamic correspondence on a null hypersurface: inconsistencies and advancement

### v1

@@ 4 Hydrodynamic description on a null surface: A formulation based on EM tensor
which, upon using our identification 40 (along with $\eta=1/16\pi$, $\zeta=-1/16\pi$, $\Theta=\theta$, and $P=\kappa/8\pi$) takes the form
@@
In the above 46, the discrepancy between Lie-derivative and convective derivative is resolved and 46 looks exactly like 8. Also note that in 46, the bulk viscosity $\Theta$ is zero. However, its derivative is non-zero. Again, let us emphasize the fact that here we have obtained the known DNS equation, which is analogous to the momentum conservation relation, in a conventional manner i.e. by taking the appropriate projection of the conservation of energy-momentum tensor. Moreover, we show that the fluid parameters can be naturally identified from the different components of the energy-momentum tensor.
Let us now dive deeper to seek an analogous relation of the energy conservation relation (i.e. analogous relation of 7) in null hydrodynamics, which does not appear in literature. For that purpose, we again follow the conventional path and contract 41 with the null generator $l^{a}$, which yields
$l^{c}\mathcal{D}_{a}T^{a}_{~{}~{}c}|_{(null)}=\frac{1}{8\pi}R_{ab}l^{a}l^{b}~{}.$ (47)
The above equation is analogous to the null Raychowdhuri equation (NRE) [25, 49]. It has been used to prove the area theorem of black holes and therefore played a pivotal role in the thermodynamic description of horizons. However, here we will show that it also has a fluid interpretation. Since in the absence of matter, the right-hand side vanishes and hence it is then analogous to 2. Therefore we expect that the null Raychoudhury equation can be interpreted as energy conservation relation in the context of fluid description of gravity. This is more vivid if one uses the adapted coordinates, described around 16.
In absence of matter, identifying $l^{a}=(1,0,\mathbf{V})$, and expanding the left hand side of 47 one obtains (for details, see C),
@@
This exactly resembles to 7 with the vanishing heat flux. Thus, we not only obtain the energy conservation relation for null hydrodynamics (which was missing earlier in the literature), we obtain it conventionally. This also suggests that our identification of null BY-like tensor as the analogous of the EMT is appropriate and the entire fluid-gravity description can be extracted from the Brown-York-like energy-momentum tensor of the null surface. One only has to take the conventional projection of 41 to obtain the desired relations. Thus this provides a conventional description of the fluid interpretation of gravity. As mentioned earlier, the choice of $\partial_{t}q_{AB}=0$ implies that in an adapted coordinate system concerning $\mathcal{H}$, the induced metric on $\mathcal{B}^{(null)}$ becomes stationary. It is precisely due to this constraint as shown in (24) that the second fundamental form becomes equivalent to the stress tensor of a viscous two-dimensional null fluid. Moreover, note that the obtained energy- and momentum conservation relations on the null surface, i.e. 49 and 44 (or 46 in a local inertial frame) matches with 7 and 8 of the fluid equations. Those fluid equations (i.e. 7 and 8) are originally derived from 5 by plugging the continuity equation with the energy- and momentum conservation relation. This implies that continuity relation in null hydrodynamics is embedded in 49 and 44 (or 46), which we obtain in the following indirect manner.
The above discussion explores the fact that the gravitational dynamics, reflected through the conservation of BY-like tensor on the null surface, can be interpreted as NS fluid under a few restrictions on the gravity side. With this setup, gravity can be interpreted as a non-relativistic viscous fluid, and the fluid parameters can be determined in terms of gravitational parameters. We already mentioned a few of them. We will see if more such parameters can be further identified. For example, comparison of 48 with 7 leads us to the identification of the mass-energy density of the effective fluid, which is given by
$\rho\epsilon^{N}=-\frac{\theta}{8\pi}~{},$ (50)
where $\rho$ is the analogue of mass density (in this case it is mass per unit area of $\mathcal{B}^{(null)}$). The momentum density is $\rho v_{A}=-\Omega_{A}/8\pi$. Transacting this relation with the fluid velocity $V^{A}$, we obtain,
@@
Thus, the continuity equation in null hydrodynamics can be framed in the form of the first equation in 5 as
@@
Comparing 51 with 50, we obtain the non-relativistic version of the internal energy density of the fluid being,
$\epsilon^{N}=-\frac{\theta}{8\pi\rho}=\frac{\theta\bm{V}^{2}}{\bm{\Omega}\cdot\bm{V}}~{}.$ (53)
@@
then using 55, 51 and the fact that $P=\kappa/8\pi$ in 54, we obtain that the entropy per unit mass of the null fluid is,
@@
where $\sqrt{q}$ is the area element on the null surface. This implies that the entropy of the null fluid per unit mass per unit area of the null surface is the usual Bekenstein Hawking entropy density. This shows that the null fluid describing the gravitational dynamics of $\mathcal{H}$ is consistent with the KSS bound with $\eta/s=1/4\pi$. The imposition on $\bm{\Omega}\cdot\bm{V}<0$, further implies via 55 and 53, that the chemical potential and the energy density of the effective fluid is negative unless $\theta<0$. However, the fact that the internal energy of gravity is negative is because of its self-attractive nature. This means that the expansion of $\mathcal{H}$ is positive, which is indeed supplemented by its teleological boundary conditions. The fact that for the null fluid, we have the temperature $T=\kappa/2\pi$ and the pressure $P=\kappa/8\pi$, leads us to the equation of state,
@@
5 Connections with previous results: NS VS Carroll fluid
The central theme in this work has been to elucidate and explore further the connections between null surface dynamics with that of a two-dimensional non-relativistic viscous fluid. In this regard, our work follows the membrane paradigm approach introduced by Price, Thorne, and MacDonald [44, 45] originating from the work of Damour [23, 24]. Along the same lines as the works done under the membrane paradigm, we also here make a $3+1$ spacetime split thus rendering a preferred choice of a time coordinate. With such a spacetime split and a given choice of adapted coordinates concerning the null surface, we essentially break away from general covariance, while analyzing the field equations on the null surface. There arise certain conceptual issues under such an analysis. The first one is regarding the bulk viscosity of the two-dimensional null fluid being negative, which is unlike ordinary fluids. This means that such a fluid is unstable under perturbations. This is related to the fact that a global null hypersurface tends to expand or contract [50]. The NRE, under the adapted coordinate system, is identified with that of the energy conservation equation of a viscous fluid (49) with no heat flux. The energy density is proportional to the expansion scalar and the NRE is a nonlinear first-order differential equation for the expansion scalar. Since the energy density is negative due to the self-attractive nature of Einstein’s gravity, this implies that the expansion scalar is positive implying that the null surface only expands under evolution along the null generators or time (under physically motivated energy conditions). This has no counterpart in ordinary fluid dynamics. To avoid this unbounded growth in the horizon area, teleological boundary conditions have to be imposed [69, 70]. That is, instead of specifying the initial conditions, one must provide the final conditions (say, for example, the horizon at the end of time-evolution reaches equilibrium defined by vanishing expansion). The growth of the null surface is acausal.
It has to be pointed out that similar counterparts have previously been explored, albeit under a different geometrical construction. That particular construction is essentially that of a stretched horizon. Now, under this purview, there have been two branches of analysis leading to a fluid description of the null surface. The first one is the membrane paradigm approach. The second one is the construction of Carrollian stretched horizons. As can be evident, these two constructions rely on the existence of a family of timelike stretched surfaces/horizons, with the null hypersurface being a limiting case in this family of stretched horizons. That is, geometrically the neighborhood of the null surface is foliated by a family of timelike hypersurfaces (as opposed to null surfaces as in our case).
In the Carrollian membrane paradigm, the vicinity of the null boundary $\mathcal{H}$ (located at a finite distance) is foliated by a family of three-dimensional timelike surfaces $\mathcal{T}$. Such a family of timelike surfaces is defined by the scalar function $r(x^{a})=\text{constant}>0$. The null boundary is obtained in the limit $r(x^{a})=0$. That is, essentially, the null limit to $\mathcal{H}$ is obtained by taking the limit $r\rightarrow 0$. However, such a family $\mathcal{T}$ is not foliated by spacelike hypersurfaces $\Sigma_{t}$ (which would in turn provide the notion of a time evolution vector field $\bm{t}$ and hence predispose the structures $\mathcal{H}$ and $\mathcal{T}$ with a Galilean picture). Instead, Carroll structures are imposed on these timelike surfaces $\mathcal{T}$ as well as the null boundary $\mathcal{H}$ [37, 38]. The normal form to these Carroll horizons $\mathcal{T}$ is given by
$\underline{\bm{s}}=e^{\bar{\alpha}}\underline{\bm{d}}r~{},$ (59)
with $\tilde{\rho}$ being a smooth scalar in $(\mathcal{M},\textbf{g},\bm{\nabla})$. The Carroll structure becomes null only on the null boundary obtained by taking $\tilde{\rho}=0$. In that case, the normal vector $\bm{s}$ coincides with the null generators $\bm{l}$ of $\mathcal{H}$. With such a Carrollian structure implemented on $\mathcal{T}$, an appropriate Carrollian fluid energy-momentum tensor can be described on the stretched horizon [37, 38]. In the null limit, this Carrollian (surface) energy-momentum tensor matches with the one introduced in [42] and (39) up to an overall negative sign. Imposing the condition that $\tilde{\rho}$ is fixed on the stretched membrane $\mathcal{T}$, the (vacuum) Einstein field equations are then shown to be conservation laws of such a Carrollian fluid energy-momentum tensor. The corresponding viscous stress tensor, pressure, energy density, momentum density, and heat current have been derived for such a Carrollian fluid [37, 38].
At this point, it becomes quite imperative to discuss the differences in the approach of stretched membranes and those employed in our study. At the very outset, the obvious constructional difference is regarding the foliation of $(\mathcal{M},\textbf{g},\bm{\nabla})$ about the null boundary $\mathcal{H}$. Both the Carrollian and the membrane paradigms employ foliating the spacetime in the vicinity of $\mathcal{H}$ by timelike hypersurfaces. In our work, following Carter, we impose a null foliation of the spacetime. By this, we avoid the issues of singularity that arise in the membrane paradigm picture by taking the null limit. We, however, in line with the membrane paradigm do employ a $3+1$ decomposition of the null boundary $\mathcal{H}$. This gives rise to a preferred time evolution vector field $\bm{t}$ take breaks general covariance bringing into the structure of $\mathcal{H}$, the Galilean picture. This is avoided in the Carrollian description that works in the ultra-relativistic regime. Another technical difference that arises in the constructions is that of the nature of the null rigging vector $\bm{k}$. For the Carrollian hydrodynamics [37, 38] on $\mathcal{T}$, the null rigging vector is assumed to generate null geodesics i.e they satisfy the auto-parallel equation ($\nabla_{\bm{k}}\bm{k}=\tilde{\kappa}\bm{k}$, $\tilde{\kappa}$ being the non-affinity parameter). This is also the same case as employed in the construction of Gaussian null coordinates w.r.t. the null boundary $\mathcal{H}$. However, in our case of the null foliation of spacetime, the auxiliary null vector field is not constrained to be geodesic.
@@ 5 Connections with previous results: NS VS Carroll fluid
One final comment we make is on the nature of a constraint that is employed to bring in the fluid description. In the case of Carrollian hydrodynamics, while considering the surface energy-momentum tensor, the scalar field $\tilde{\rho}$ (describing the non-nullness of the normal one form to $\mathcal{T}$) is assumed to be constant over the Carrollian membrane $\mathcal{T}$. This can be suitably done for a choice of the scalar function $\bar{\alpha}$. In our case, we have employed the constraint, that the induced metric on the transverse cut $\mathcal{B}^{(null)}$ of $\mathcal{H}$ is stationary and hence independent of the time parameter. This can be done with a preferred choice of an adapted coordinate system w.r.t $\mathcal{H}$.
@@ 6 Discussions
The curious relation found by Damour [23, 24] in 1979 (i.e. 27) tempted physicists to draw the analogy between fluid dynamics with the dynamics of the null surface. However, there are no other connections in support of the connection. Furthermore, the DNS equation is not exactly the same as the NS equation as we discussed earlier. Hence, it made physicists skeptical regarding the potential connection between fluid dynamics and the dynamics of the null surface. For example, Price and Throne do not connect Damour’s result with the NS equation [36]. Instead, they refer to it as the Hajicek equation. To resolve the difference between the DNS equation and NS equation, Padmanabhan [26] suggested a local inertial frame, in which case the DNS equation and NS equation become identical. However, there are no physical reasons for preferring such a frame. Furthermore, since there are no other relations that show the connection between fluid dynamics and the dynamics of the null surface, it was also not possible to examine whether the identifications of fluid parameters in DNS equations are consistent. In addition, the quantity that might play the role of EMT in null hydrodynamics was also unknown.
@@
To address these issues, we introduce our analysis. Our objective has been to explore and strengthen the study of the emergent paradigm of gravity, specifically from the hydrodynamic setting. Observing that in this body of work, there does not exist a complete picture, we delved ourselves into this venture to provide some of the missing links. The central theme in this work is interpreting the conservation of a well-defined null Brown-York tensor on $\mathcal{H}$ as providing a hydrodynamical description of the null surface dynamics as that of a two-dimensional null viscous fluid. Knowing that the true dynamical degrees of freedom of the (degenerate) null surface $\mathcal{H}$ being $\mathcal{P}_{ab}$ and $\mathcal{P}^{(l)}_{a}$, which are the conjugate variables to $q^{ab}$ and $l^{a}$, we explored and showed that the gravitational dynamics of $\mathcal{H}$ can be encoded in terms of these quantities. Using a well-defined null BY tensor defined on $\mathcal{H}$ which plays the role of the EMT, we have shown how all the central hydrodynamical equations can be derived as (vacuum) conservation laws of such a null BY tensor. For such a null BY tensor, the corresponding gravitational counterparts of energy density, momentum density, and spatial stress tensor can be well defined. If the null surface dynamics is indeed to be interpreted as a null fluid, then these gravitational counter-parts (of energy density, momentum density, and spatial stress) should also be extracted from the fluid dynamical equations as well. This is precisely what we did as consistency checks by interpreting the null surface dynamics as conservation laws of the null BY tensor (in the vacuum case). With the conservation law of the null BY tensor projected onto the spatial cross section $\mathcal{B}^{(null)}$, we obtained the DNS equation, which is equivalent to the momentum conservation law in null hydrodynamics. Similarly, projecting the conservation equation of the null BY tensor along the null generators of $\mathcal{H}$ and exploring the resulting dynamics in a stationary adapted coordinate system, we obtained the energy conservation law of null hydrodynamics. We notice that the energy density of such a null fluid is negative and its heat current vanishes. Our analysis shows that the suitable projections of the conservation law of the null BY tensor give rise to the necessary hydrodynamical laws of the null fluid. For such a null fluid and suitable identifications of temperature and chemical potential, we used the (densitized) Euler relation to indeed verify that the entropy density in the vacuum Einstein case is the usual Bekenstein-Hawking entropy density and that this null fluid satisfies the KSS bound.
Let us again reiterate that such analogous interpretations have been explored previously. Especially, in the membrane paradigm and Carrollian membrane paradigm, similar interpretations have been previously brought forward. We, however, have worked in a geometric setup that is distinct from the stretched (timelike) case. As opposed to the Carrolian membrane paradigm, where it has been suggested that the symmetries of the null surface $\mathcal{H}$ are Carroll symmetries, we however have analyzed the null surface dynamics in a $3+1$ setting choosing a preferred time evolution vector field and adapted choice of coordinates. Thus the fluid setting that we deal with is the usual Galilean symmetries that are evident in the body of work pertaining to the membrane paradigm and emergent gravity paradigm. However, owing to the null foliation of the spacetime, we do not encounter the possible singularities that arise on taking the null limit of stretched membranes in the membrane paradigm. We have hopefully strengthened the picture of emergent gravity by providing a consistent viewpoint of deriving all the necessary hydrodynamical content of the null surface dynamics arising from the conservation law of a well-defined null Brown York tensor.

### v2

@@ 5 Is there any non-relativistic counterpart?
which, upon using our identification Eq. ( (along with $\eta=1/16\pi$, $\zeta=-1/16\pi$, $\theta=\theta$, and $P=\kappa/8\pi$) takes the form
@@
In the above Eq. (, the discrepancy between Lie-derivative and convective derivative is resolved and Eq. ( looks exactly like Eq. (. Again, let us emphasize the fact that here we have obtained the known DNS equation, which is analogous to the momentum conservation relation, in a conventional manner i.e. by taking the appropriate projection of the conservation of energy-momentum tensor. Moreover, we show that the fluid parameters can be naturally identified from the different components of the energy-momentum tensor.
As promised in the previous section, we now demonstrate that Eq. (, the null Raychaudhuri equation, obtained as a particular projection of the conservation of the null Brown-York tensor, admits a fluid–dynamic interpretation when expressed in adapted coordinates. In absence of matter, identifying $l^{a}=(1,0,\mathbf{V})$, and expanding the left hand side of Eq. ( one obtains (for details, see Section D),
@@
This exactly resembles to Eq. ( with the vanishing heat flux. Thus, we not only obtain the energy conservation relation for null hydrodynamics (which was missing earlier in the literature), we obtain it conventionally. This also suggests that our identification of null BY-like tensor as the analogous of the EMT is appropriate and the entire fluid-gravity description can be extracted from the Brown-York-like energy-momentum tensor of the null surface. One only has to take the conventional projection of Eq. ( to obtain the desired relations. Thus this provides a conventional description of the fluid interpretation of gravity. As mentioned earlier, the choice of $\partial_{t}q_{AB}=0$ implies that in an adapted coordinate system concerning $\mathcal{H}$, the induced metric on $\mathcal{B}^{\rm(null)}$ becomes stationary. It is precisely due to this constraint as shown in (Eq. () that the second fundamental form becomes equivalent to the stress tensor of a viscous two-dimensional null fluid. Moreover, note that the obtained energy- and momentum conservation relations on the null surface, i.e. Eq. ( and Eq. ( (or Eq. ( in a local inertial frame) matches with Eq. ( and Eq. ( of the fluid equations. Those fluid equations (i.e. Eq. ( and Eq. () are originally derived from Eq. ( by plugging the continuity equation with the energy- and momentum conservation relation. This implies that continuity relation in null hydrodynamics is embedded in Eq. ( and Eq. ( (or Eq. (), which we obtain in the following indirect manner.
The above discussion explores the fact that the gravitational dynamics, reflected through the conservation of BY-like tensor on the null surface, can be interpreted as NS fluid under a few restrictions on the gravity side. With this setup, gravity can be interpreted as a non-relativistic viscous fluid, and the fluid parameters can be determined in terms of gravitational parameters. We already mentioned a few of them. We will see if more such parameters can be further identified. For example, comparison of Eq. ( with Eq. ( leads us to the identification of the mass-energy density of the effective fluid, which is given by
$\rho\epsilon^{\rm N}=-\frac{\theta}{8\pi}~,$ (51)
where $\rho$ is the analogue of mass density (in this case it is mass per unit area of $\mathcal{B}^{\rm(null)}$). The momentum density is $\rho v_{A}=-\Omega_{A}/8\pi$. Transacting this relation with the fluid velocity $V^{A}$, we obtain,
@@
Thus, the continuity equation in null hydrodynamics can be framed in the form of the first equation in Eq. ( as
@@
Comparing Eq. ( with Eq. (, we obtain the non-relativistic version of the internal energy density of the fluid being,
$\epsilon^{\rm N}=-\frac{\theta}{8\pi\rho}=\frac{\theta\bm{V}^{2}}{\bm{\Omega}\cdot\bm{V}}~.$ (54)
@@
then using Eq. (, Eq. ( and the fact that $P=\kappa/8\pi$ in Eq. (, we obtain that the entropy per unit mass of the null fluid is,
@@
where $\sqrt{q}$ is the area element on the null surface. This implies that the entropy of the null fluid per unit mass per unit area of the null surface is the usual Bekenstein Hawking entropy density. This shows that the null fluid describing the gravitational dynamics of $\mathcal{H}$ is consistent with the KSS bound with $\eta/s=1/4\pi$. The imposition on $\bm{\Omega}\cdot\bm{V}<0$, further implies via Eq. ( and Eq. (, that the chemical potential and the energy density of the effective fluid is negative unless $\theta<0$. However, the fact that the internal energy of gravity is negative is because of its self-attractive nature. This means that the expansion of $\mathcal{H}$ is positive, which is indeed supplemented by its teleological boundary conditions. The fact that for the null fluid, we have the temperature $T=\kappa/2\pi$ and the pressure $P=\kappa/8\pi$, leads us to the equation of state,
@@
One final comment we make is on the nature of a constraint that is employed to bring in the fluid description. In the case of Carrollian hydrodynamics, while considering the surface energy-momentum tensor, the scalar field $\tilde{\rho}$ (describing the non-nullness of the normal one form to $\mathcal{T}$) is assumed to be constant over the Carrollian membrane $\mathcal{T}$. This can be suitably done for a choice of the scalar function $\bar{\alpha}$. In our case, we have employed the constraint, that the induced metric on the transverse cut $\mathcal{B}^{\rm(null)}$ of $\mathcal{H}$ is stationary and hence independent of the time parameter. This can be done with a preferred choice of an adapted coordinate system w.r.t $\mathcal{H}$.
Before going to the discussion section let us conclude this section with the following comments. The membrane paradigm approach, introduced by Price, Thorne, and MacDonald [69, 55], was originated from the work of Damour [23, 24]. Along the same lines as the works done under the membrane paradigm, we also here make a $3+1$ spacetime split thus rendering a preferred choice of a time coordinate. With such a spacetime split and a given choice of adapted coordinates concerning the null surface, we essentially break away from general covariance, while analyzing the field equations on the null surface. There arise certain conceptual issues under such an analysis. The first one is regarding the bulk viscosity of the two-dimensional null fluid being negative, which is unlike ordinary fluids. This means that such a fluid is unstable under perturbations. This is related to the fact that a global null hypersurface tends to expand or contract [78]. The NRE, under the adapted coordinate system, is identified with that of the energy conservation equation of a viscous fluid (Eq. () with no heat flux. The energy density is proportional to the expansion scalar and the NRE is a nonlinear first-order differential equation for the expansion scalar. Since the energy density is negative due to the self-attractive nature of Einstein’s gravity, this implies that the expansion scalar is positive implying that the null surface only expands under evolution along the null generators or time (under physically motivated energy conditions). This has no counterpart in ordinary fluid dynamics. To avoid this unbounded growth in the horizon area, teleological boundary conditions have to be imposed [71, 86]. That is, instead of specifying the initial conditions, one must provide the final conditions (say, for example, the horizon at the end of time-evolution reaches equilibrium defined by vanishing expansion). The growth of the null surface is acausal.
@@ 6 Discussions
The curious relation found by Damour [23, 24] (i.e. Eq. () within null foliation formalism tempted physicists to draw the analogy between fluid dynamics with the dynamics of the null surface. However, the DNS equation is not exactly the same as the NS equation as we discussed earlier. Hence, it made physicists skeptical regarding the potential connection between fluid dynamics and the gravitational dynamics on the null surface. For example, Price and Throne do not connect Damour’s result with the NS equation [33]. Instead, they refer to it as the Hajicek equation. To resolve the difference between the DNS equation and NS equation, Padmanabhan [26] suggested a local inertial frame, in which case the DNS equation and NS equation become identical. However, there are no physical reasons for preferring such a frame. Furthermore, since there are no other relations that show the connection between fluid dynamics and the dynamics of the null surface, it was also not possible to examine whether the identifications of fluid parameters in DNS equations are consistent. In addition, the quantity that might play the role of EMT in null hydrodynamics was also unknown.
@@
To address these issues, we introduce our analysis. Our objective has been to explore and strengthen the study of the emergent paradigm of gravity, specifically from the hydrodynamic setting. We reiterate that previous works have indeed illuminated upon the connections with gravitational dynamics and hydrodynamics on an internal null boundary. Chief among them have been the Membrane paradigm and Carroll hydrodynamics. Both these works insist on studying and analyzing the gravitational dynamics of an internal null hypersurface in terms of appropriate fluid dynamics and their symmetries. Central to both these formalisms have been the idea of studying the gravitational dynamics on the internal null boundary $\mathcal{H}$ by foliating the spacetime in the vicinity of $\mathcal{H}$ by a family of timelike hypersurfaces with an appropriate way of taking the null limit to $\mathcal{H}$. The key difference in this work has been essentially to study the gravitational dynamics of $\mathcal{H}$ by foliating the neighborhood of $\mathcal{H}$ by a family of null surfaces. This is mathematically pleasing as there is no need to take a null limit and their are no issues of divergences/singularities that can occur. In tandem with the previous works of the Membrane and Carroll membrane paradigms, we delved ourselves into this venture of filling in the procedures in this set-up where the spacetime in the vicinity of the internal null boundary is foliated by a null family. For this, in such a null foliated construction, we proceeded to use the appropriate null energy momentum tensor. The central theme then is interpreting the conservation of such a well-defined null Brown-York tensor on $\mathcal{H}$ as providing a hydrodynamical description of the null surface dynamics as that of a two-dimensional null viscous fluid. Knowing that the true dynamical degrees of freedom of the (degenerate) null surface $\mathcal{H}$ being $\mathcal{P}_{ab}$ and $\mathcal{P}^{(l)}_{a}$, which are the conjugate variables to $q^{ab}$ and $l^{a}$, we explored and showed that the gravitational dynamics of $\mathcal{H}$ can be encoded in terms of these quantities. Using the well-defined null BY tensor defined on $\mathcal{H}$ which plays the role of the EMT, we have shown how all the central hydrodynamical equations can be derived as (vacuum) conservation laws of such a null BY tensor. For such a null BY tensor, the corresponding gravitational counterparts of energy density, momentum density, and spatial stress tensor can be well defined. If the null surface dynamics is indeed to be interpreted as a null fluid, then these gravitational counter-parts (of energy density, momentum density, and spatial stress) should also be extracted from the fluid dynamical equations as well. This is precisely what we did as consistency checks by interpreting the null surface dynamics as conservation laws of the null BY tensor (in the vacuum case). With the conservation law of the null BY tensor projected onto the spatial cross section $\mathcal{B}^{\rm(null)}$, we obtained the DNS equation, which is equivalent to the momentum conservation law in null hydrodynamics. Similarly, projecting the conservation equation of the null BY tensor along the null generators of $\mathcal{H}$ and exploring the resulting dynamics in a stationary adapted coordinate system, we obtained the energy conservation law of null hydrodynamics. We notice that the energy density of such a null fluid is negative and its heat current vanishes. Our analysis shows that the suitable projections of the conservation law of the null BY tensor give rise to the necessary hydrodynamical laws of the null fluid. For such a null fluid and suitable identifications of temperature and chemical potential, we used the (densitized) Euler relation to indeed verify that the entropy density in the vacuum Einstein case is the usual Bekenstein-Hawking entropy density and that this null fluid satisfies the KSS bound.
We have hopefully strengthened the picture of emergent gravity by providing a consistent viewpoint of deriving all the necessary hydrodynamical content of the null surface dynamics arising from the conservation law of a well-defined null Brown York tensor. Hence we hope that with this work, the absence of fluid description of gravity on a generic null surface by adopting null foliation in the previous literature, has now been filled.
A natural question that may arise in this context is what happens when we perform the analysis on null infinity rather than an internal null boundary at finite distances. In this regard, we became aware of a work [87] dealing precisely with this question, which came after we submitted our manuscript for peer review. In the work, similar to our geometrical construction, the author considered a unified framework by foliating the bulk spacetime with a family of null surfaces and consequently studying the null surfaces within the bulk in the Bondi-Sachs gauge and then sending off this null surface towards future null infinity. Using the Carrollian data on the internal surface, they study the projections of the Einstein’s equations (specifically the null Raychaudhuri and the Damour Navier-Stokes ones) on the null surface and show that at null infinity they asymptote to the Einstein’s equation for the boundary metric. By assuming that the boundary metric is time independent (which is analogous to our constraint $\partial_{t}q_{AB}=0$), the author quite interestingly shows that the asymptotic limit of the time component of the conservation of the null Brown-York stress tensor (which effectively yields the null Raychaudhuri equation) leads to the Bondi mass-loss formula [88]. The author provides a hint that asymptotic limit of the Damour Navier-Stokes equation must be related to the angular momentum equation involving the angular momentum aspect [88]. The author also proposes the asymptotic limit of the null Brown-York stress tensor and shows that the phase space of the internal null boundary asymptotes to the Ashtekar-Struebel phase space [89] of null infinity.